OpenStudy (anonymous):

really need help with systems of equationss!! 2x-4y=12 3x-6y=21 (I need to find x and y)

OpenStudy (e.mccormick):

Know how to do elimination or substitution?

OpenStudy (anonymous):

2x – 4y = 12.....A 3x – 6y = 21.....B from equation A find the value of x or y its up to your choice and then place it in equation B in this way you will get the answer of A and B

OpenStudy (anonymous):

I know I need to find one value first but I do not know how to get that first value

OpenStudy (anonymous):

that is exactly what I am learning but I dont quite get it

OpenStudy (e.mccormick):

Well, with this one there is something you can do first that will help. Multiply the entire first equation by 1/2 and the 2nd by 1/3rd. You should see something significant then.

OpenStudy (anonymous):

2x-4y=12 place x on one side and rest of values on other side lie this 2x=12+4y x=12+4y/2 x=12/2+4y/2 x=6+2y .....in this way you get the value of x now put this value in second equationto get the value of y like this 3x-6y=21 3(6+2y)-6y=21 18+8y-6y=21 8y-6y=21-18 2y=3 y=3/2 so in this way you get the value of y as well x=6+2y x=6+2(3/2) x=6+3 x=9

OpenStudy (anonymous):

THANK YOUU so muchhh!!!

OpenStudy (anonymous):

welcome

OpenStudy (e.mccormick):

Interesting idea, except it is wrong. That is not the answer.

OpenStudy (e.mccormick):

You made your mistake here: 3(6+2y)-6y=21 18+8y-6y=21

OpenStudy (anonymous):

oh one mistake

OpenStudy (anonymous):

@epoweritheta is this right

OpenStudy (e.mccormick):

Now, if you correct that, or do what I said at the start, or graph it, you will see something significant.

OpenStudy (anonymous):

no solution right?

OpenStudy (e.mccormick):

\(2x-4y=12 \) \(3x-6y=21\) \(\dfrac{1}{2}(2x-4y)=\dfrac{1}{2}(12) \) \(\dfrac{1}{3}(3x-6y)=\dfrac{1}{3}(21)\) \(x-2y=6\) \(x-2y=7\) See thee issue now? That means the same as saying \(6=7\).

OpenStudy (e.mccormick):

Yes. They are parallel lines, so no solution: https://www.desmos.com/calculator/8ikbblrbte

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

could you help me with another one please?

OpenStudy (e.mccormick):

Sure, just tag me when you open the new question.

OpenStudy (anonymous):

great thankyouu!